# FindingDory: A Benchmark to Evaluate Memory in Embodied Agents

### v1

@@ 4 Experiments
We evaluate several baseline approaches on FindingDory and assess their ability to complete memory-intensive tasks using a combination of vision-language reasoning and navigation. Below, we describe the baseline architecture, implementation variants, and evaluation metrics.
Figure 2: FindingDory Task Performance: Proprietary VLMs such as Gemini-2.0-flash and GPT-4o fail to perform appreciably across most FindingDory task categories. Agents struggle significantly on multi-goal tasks where multiple subgoals are to be achieved sequentially based on the interaction videos collected during Phase 1 of the task (sometimes in a specific order). The Qwen-SFT baseline (see Sec. D.5) which is trained to predict ground-truth frame indices performs best but still reaches only $\approx 50\%$ on the high-level tasks. This highlights a considerable gap in the spatio-temporal reasoning capabilities of frontier models.
@@ 4.1 Evaluated Approaches
To establish strong baselines, we adopt a hierarchical architecture (Fig. 11) comprising two modules:
(1) a high-level goal selection module, which processes the interaction history to select goal frames, and
(2) a low-level navigation policy, which executes actions to reach the selected frame.
This architecture enables reasoning over long-horizon multimodal experience while delegating fine-grained control to a learned policy. Inspired by prior work [45], we avoid directly predicting actions using VLMs due to their poor performance in continuous control, and instead restrict them to high-level goal prediction.
@@ 4.1.1 High-Level Goal Selection
Video LM Agent. This agent receives the full interaction video along with a task instruction. Each frame in the video is annotated with its index and the time of day at which it was captured. The agent is tasked with predicting the index of a frame from the video that corresponds to a viable goal state. Specifically, the predicted frame should represent a viewpoint from which, if the agent were to navigate back to the same camera position, it could successfully complete the task. Although the model outputs a single frame index, selecting the correct one often requires reasoning over multiple temporally scattered observations, including frames where the agent may not be actively interacting. Since different VLMs are capable of handling different numbers of frames in context, we subsample the frames passed to each model based on what empirically performs best. We evaluate both proprietary (Gemini-2.0-Flash [39], GPT-4o) and open-source (Qwen2.5-VL [5], Gemma-3 [40]) vision-language models capable of processing long video contexts. See Secs. D.1, D.2 and D.3 for more details.
Textual Memory Agent. Similar to [25, 4], we also consider an explicit, memory-building baseline which breaks down the video frames into chunks and uses a VLM to generate textual summaries for each video chunk. After “captioning” the entire video sequence, the VLM selects frame indices by reasoning over the generated textual descriptions. See Sec. D.4 for more details.
Supervised Fine-Tuning (SFT). We additionally explore a supervised fine-tuning approach where the high-level VLM is trained to predict goal frame indices, given the task prompt and interaction history. Since multiple frames may validly represent the goal (e.g., several consecutive frames showing an object being placed), we supervise the model on the full list of acceptable target frames. At test time, a single frame is sampled from the list of predictions and is considered correct if it matches any of these valid frames. This setup enables the model to learn task-specific cues from examples and serves as a stronger baseline than zero-shot prompting, particularly for tasks requiring subtle temporal grounding. We provide additional details in Sec. D.5.
@@ 4.1.2 Low-Level Navigation
To execute the high-level plan, we use a low-level policy that takes the selected goal frame and outputs discrete actions: MOVE-FORWARD, TURN-RIGHT, TURN-LEFT, and STOP. Our primary policy is the LSTM-based controller from OVRLv2 [46], trained to reach image goals from egocentric RGB-D observations (see Appendix E.1).
We also evaluate a mapping-based policy that deterministically navigates to the location corresponding to the goal frame using a global map, without relying on learned visual features (Appendix E.2).
Solvability. Since the action space of our high-level baseline is constrained to frames observed during the experience collection phase, there are tasks for which this is insufficient. For example, in the task “Navigate to a room that you did not visit yesterday,” all available waypoints correspond to previously visited locations, making it impossible to satisfy the instruction regardless of the selected frame. To establish an upper bound on task difficulty, we quantify the percentage of episodes that could theoretically be solved through optimal goal selection from interaction history, assuming perfect navigation (teleportation) to the selected waypoint. This upper bound is the dashed line (Oracle) in Fig. 2.
(a) Relaxed metrics for Objects.
(b) Relaxed metrics for Receptacles.
(c) HL-SR vs HL-SPL Comparison.
(d) Hierarchical policy results.
Figure 3: (a, b) Relaxed metrics for objects and receptacles tasks. We find that the models significantly improve in relaxed metrics for tasks involving receptacles while tasks targeting smaller objects do not exhibit similar improvements. (c) Success Rate vs SPL performance On tasks in which the VLM has the option to choose between multiple correct frames, we find that all models choose suboptimal frames sometimes resulting in the reduced HL-SPL metric. (d) Hierarchical policy results on FindingDory When Qwen is combined with an ImageNav or Mapping agent, the overall SR and SPL reduces significantly (seen in the LL-SR/LL-SPL metrics).
@@ 4.2 Evaluation Metrics
We benchmark agent performance on the FindingDory tasks using Low Level Success Rate (LL-SR), assessing whether the agent finds the right target entities, and Low Level Success-weighted-by-Path-Length (LL-SPL), evaluating if the agent selects and navigates to the optimal entity using the shortest possible path. Since we use a hierarchical baseline, we additionally introduce a High-Level Policy Success Rate (HL-SR), and High-Level Policy SPL (HL-SPL) which measures the accuracy and efficiency of the high-level policy in selecting the correct and closest goal frames for navigation. To further analyze failure modes, we also report two relaxed variants: Distance-to-Goal-Only SR (DTG-SR), which relaxes the requirement that the target object must be visible in the selected frame, and Semantic Coverage SR (SC-SR), which ignores spatial proximity. Additional metrics, success criteria, and threshold specifics are provided in Sec. B.3.
@@ 5 Results
In this section, we analyze the performance of the outlined baselines on FindingDory, presenting results grouped according to the categories described in Section 3.2.
@@
State-of-the-Art VLMs struggle on the FindingDory benchmark. From  Fig. 2, we observe consistently low success rates on all task subsets across baselines, with GPT-4o achieving the highest performance at 27.3%, Gemini-2.0-Flash reaching 25.7%, while Qwen2.5-VL and Gemma-3 score in the 13–16% range. The text-based agent performs the worst overall, underscoring that textual memory representations alone are insufficient for disambiguating visual goals or reasoning over complex visual histories. Notably, the supervised fine-tuned (SFT) variant outperforms all frozen VLMs, achieving an average improvement of 25% across tasks, with considerable improvement on temporal and multi-goal tasks.
@@
In single-goal tasks, all methods struggle particularly with the spatial relationship subset, with even GPT-4o achieving only 14.5% success. These tasks require reasoning about relative distances between entities and the agent—an ongoing challenge noted in prior work [47]. In contrast, room identification tasks yield the highest performance, likely due to the small number of distinct rooms and the lenient goal condition of reaching any location within the correct room. Performance on conditional interaction and object attribute tasks is comparable, suggesting that current VLMs handle both category-based queries (e.g., “apple”) and attribute-based descriptions (e.g., “red food”) with similar effectiveness.
@@
Temporal tasks pose additional challenges to the existing VLMs, with models consistently failing to identify the correct order of interactions—requiring multi-hop reasoning. Duration tracking is similarly difficult as it requires inferring elapsed duration based on specific points in the video. In contrast, for the time-based tasks we observe that baselines perform better as these tasks only require single frame retrieval without additional reasoning.
(a) Comparison of different video lengths
(b) Reasons for ImageNav Policy Errors
@@
Poor performance on multi-goal tasks. From Fig. 2, we observe that performance of each VLM baseline that we consider is close to zero on most multi-goal tasks. This suggests that VLMs find it hard to track and remember multiple object-receptacle interactions. For the multi-goal evaluations, we use a single-shot prompt and require the high level VLM agent to predict all target frame indices in one single response - primarily to avoid multiple inferences with long sequence lengths due to large number of frames. Qualitatively, we observe that the VLM generally fails to even predict the correct number of targets required to solve the multi-goal task. Our trained Qwen model performs slightly better, achieving close to 20% accuracy, but still falls significantly short of the achievable result of 99%.
VLMs can identify target receptacles but not precisely locate them spatially. We conduct an experiment to analyze the performance differences of VLMs between tasks involving objects versus receptacle entities. We extract a subset of FindingDory tasks, categorizing them based on whether the valid entities are objects or receptacles. From  Fig. 3(b), we observe that the relaxed Semantic Coverage-SR (SC-SR) metrics (see  Section B.3 for details) significantly increases for tasks involving receptacle entities. The marked increase in high-level success rates with relaxed distance threshold (DTG-SR) suggests that VLMs are generally able to identify frames from which large entities (receptacles) can be viewed in the image but cannot accurately identify specific frames which are spatially closest to the target entity. In contrast, tasks targeting smaller objects do not exhibit the same improvement (see  Fig. 3(a)), as fewer number of frames clearly show these small objects from farther distances. This highlights a significant limitation of current VLMs in constructing effective spatial representations to accurately localize target entities from the provided observation stream. We explicitly prompt high-level VLM agents to identify frames spatially closest to target entities, essential for task success (see Figs. 6 and 8 for example prompts).
Policy decomposition causes performance degradation. To benchmark the hierarchical agent proposed in  Sec. 4.1, we experiment using Qwen2.5-VL-7B as the high-level VLM module and image-goal navigation policy and mapping policy as the low-level agents. The low-level policy activates only when the high-level VLM correctly selects the subgoal. As depicted in  Fig. 3(d), the Low-Level-SR (LL-SR) metric significantly drops across all FindingDory tasks for the image-goal navigation policy. We attribute this decline to a distribution shift in the visual goals provided to the navigation policy. During training, the policy encounters randomly sampled start and end goal poses, differing substantially from those selected by VLMs during evaluation. Successful VLM-selected frames typically correspond closely to pick place interaction routines, resulting in image goals predominantly containing background walls and partial views of receptacle surfaces. Consequently, this limited visual context hinders effective navigation. In comparison, the mapping based navigation policy performs better, having only a 25% relative drop in SR. The SPL of both the methods remain in the single digits. This emphasizes the importance of the FindingDory benchmark in revealing these gaps which motivates the development of more reliable memory and navigation strategies
@@
Frame Subsampling. We analyze how the number of input frames affects high-level performance by evaluating agents on subsampled interaction videos of varying lengths (Fig. 4(a)). This helps assess whether models benefit from longer histories or simply overfit to sparse visual cues. Interestingly, we observe that frozen VLMs—despite being capable of accepting long contexts—do not improve with more frames. In fact, their performance often degrades at higher frame counts (minimal subsampling), suggesting that they struggle to extract relevant information from densely packed, unfiltered input. This highlights a key limitation: naïvely increasing context length does not help unless the model can effectively attend to localized events.
In contrast, the fine-tuned model (Qwen SFT) demonstrates clear gains when trained with longer videos, leveraging additional context to improve reasoning. This indicates that with appropriate supervision, models can move beyond shallow matching and utilize richer temporal signals from the interaction history.
Sanity checks. We investigate how a VLM would perform at short-horizon visual comprehension on images from the same episodes used for long-horizon evaluation in our benchmark. This tests if poor performance of the evaluated models on our long-horizon reasoning tasks is due to a domain or sim2real gap. We assess this by sampling frames around object interactions, asking models to list out detected objects, and check accuracy compared to the ground truth category via string matching – a conservative metric that misses correct predictions using synonyms or alternative phrasings. This still results in image sequences of length 50-150, but Gemini achieves an accuracy of 67% which is significantly higher than it’s average high level success rate across tasks in Fig. 2. This shows that despite clear, close-up views, models struggle to reason over them when temporal sequence comprehension is required.
Low level navigation policy failures. From  Fig. 4(b), we observe that the low-level image-goal navigation policy can fail due to invoking the termination (stop) action incorrectly or a navigation routine timeout (max budget of $1000$ steps including the data collection phase). To further understand this behavior, we analyze if the navigation policy actually looks at the target entity when the stop action is invoked. Interestingly, we find that in several instances the target entity is visible from a slightly larger distance threshold (2.0m) from the chosen task success threshold (1.0 m, in case of receptacles). This often happens when VLMs select goal frames in which the target entity is visible in the 2D image but not necessarily closest in 3D space, and thus the goal frame coordinates are slightly further away than the threshold distance.

### v2

@@ 4 Experiments
We evaluate several baseline approaches on FindingDory and assess their ability to complete memory-intensive tasks using a combination of vision-language reasoning and navigation. Below, we describe the baseline architecture, implementation details, and evaluation metrics.
@@ 4.1 Evaluated Approaches
To establish strong baselines, we adopt a hierarchical architecture (fig. 15) comprising two modules: (1) a high-level goal selection module, which processes the interaction history to select a sequence of goal frames, and (2) a low-level navigation policy, which executes actions to reach the selected frame. This architecture enables reasoning over long-horizon multimodal experience while delegating fine-grained control to a learned policy. Inspired by prior work (Xu et al., 2024), we avoid directly predicting actions using VLMs due to their poor performance in continuous control, and instead restrict them to high-level goal prediction.
@@ 4.1.1 High-Level Goal Selection
Video LM Agent. This agent receives the full interaction video along with a task instruction. Each frame in the video is annotated with its index and the time of day at which it was captured. The agent is tasked with predicting the index of a frame from the video that corresponds to a viable goal state. Specifically, the predicted frame should represent a viewpoint from which, if the agent were to navigate back to the same camera position, it could successfully complete the task. Although the model outputs a single frame index, selecting the correct one often requires reasoning over multiple temporally scattered observations, including frames where the agent may not be actively interacting. Since different VLMs are capable of handling different numbers of frames in context, we subsample the frames passed to each model based on what empirically performs best. We evaluate both proprietary (Gemini-2.0-Flash (Team et al., 2024), GPT-4o) and open-source (Qwen2.5-VL (Bai et al., 2025), Gemma-3 (Team et al., 2025), GLM-4.1V-Thinking (Hong et al., 2025)) vision-language models capable of processing long video contexts. See sections D.1, D.2 and D.3 for more details.
Textual Memory Agent. Similar to (Min et al., 2024; Anwar et al., 2024), we also consider an explicit, memory-building baseline which breaks down the video frames into chunks and uses a VLM to generate relevant textual summaries for each video chunk. After this guided “captioning” of the entire video sequence, an LLM selects goal frame indices by reasoning over the generated textual descriptions. See section D.4 for more details.
Supervised Fine-Tuning (SFT). We additionally explore a supervised fine-tuning approach where the high-level VLM is trained to predict goal frame indices, given the task prompt and interaction history. Since multiple frames may validly represent the goal (e.g., several consecutive frames showing an object being placed), we supervise the model on the full list of acceptable target frames. At test time, a single frame is sampled from the list of predictions and is considered correct if it matches any of these valid frames. This setup enables the model to learn task-specific cues from examples and serves as a stronger baseline than zero-shot prompting, particularly for tasks requiring subtle temporal grounding. We provide additional details in section D.5.
@@ 4.1.2 Low-Level Navigation
To execute the high-level plan, we use a low-level policy that takes the selected goal frame and outputs discrete actions: MOVE-FORWARD, TURN-RIGHT, TURN-LEFT, and STOP. Our primary policy is the LSTM-based controller from OVRLv2 (Yadav et al., 2023), trained to reach image goals from egocentric RGB-D observations (see section E.1) (ImageNav). We also evaluate a mapping-based policy that deterministically navigates to the location corresponding to the goal frame using a global map, without relying on learned visual features (see section E.2).
Solvability. The action space of our high-level baseline is constrained to frames observed during experience collection, which is insufficient for some tasks. For example, in the task “Navigate to a room that you did not visit yesterday” all frames correspond to previously visited locations, making it impossible to complete the instruction. To establish an upper bound on task difficulty, we quantify the percentage of episodes that could theoretically be solved through optimal frame selection from interaction history, assuming perfect navigation (teleportation) to the selected waypoint. This upper bound is the dashed line (Oracle) in fig. 2.
@@ 4.2 Evaluation Metrics
We benchmark agent performance on the FindingDory tasks using Low Level Success Rate (LL-SR), assessing whether the agent finds the right target entities, and Low Level Success-weighted-by-Path-Length (LL-SPL), evaluating if the agent selects and navigates to the optimal entity using the shortest possible path. Since we use a hierarchical baseline, we additionally introduce a High-Level Policy Success Rate (HL-SR), and High-Level Policy SPL (HL-SPL) which measures the accuracy and efficiency of the high-level policy in selecting the correct and closest goal frames for navigation. To further analyze failure modes, we also report two relaxed variants: Distance-to-Goal-Only SR (DTG-SR), which relaxes the requirement that the target object must be visible in the selected frame, and Semantic Coverage SR (SC-SR), which ignores spatial proximity. Additional metrics, success criteria, and threshold specifics are provided in section B.3.
@@ 5 Results
Figure 2: FindingDory Task Performance: Proprietary VLMs such as Gemini-2.0-flash and GPT-4o show little success across most FindingDory task categories. Agents struggle especially on multi-goal tasks where multiple subgoals must be achieved sequentially from the interaction videos collected during Phase 1. The Qwen-SFT baseline (see section D.5), trained to predict ground-truth frame indices, performs best but still reaches only $\approx 50\%$ on high-level tasks.
@@
In this section, we analyze the performance of the previously outlined baselines on FindingDory tasks, presenting results grouped according to the task categories described in section 3.1.
@@
State-of-the-Art VLMs struggle on the FindingDory benchmark. We first evaluate different VLMs on the high-level task and report success rates in fig. 2, observing consistently low performance across all task subsets (see table 4 for numbers). GPT-4o achieves the highest success at 27.3%, followed by Gemini-2.0-Flash at 25.7%, and GLM-4.1V-Thinking at 23.5%. The non-reasoning open-source models Qwen2.5-VL and Gemma-3 score between 13% and 16%. The text-based agent performs the worst overall, underscoring that textual memory representations alone are insufficient for disambiguating visual goals or reasoning over complex visual histories. Notably, the supervised fine-tuned (SFT) variant outperforms all frozen VLMs, achieving an average improvement of 25% across tasks, with considerable improvement on temporal and multi-goal tasks.
@@
In single-goal tasks, all methods struggle particularly with the spatial relationship subset, with GPT-4o achieving only 14.5% success. These tasks require reasoning about relative distances between entities and the agent, an ongoing challenge noted in prior work (Yang et al., 2024). In contrast, room identification tasks yield the highest performance, likely due to the small number of distinct rooms and the lenient success criteria of reaching any location within the correct room. Performance on conditional interaction and object attribute tasks is comparable, suggesting that current VLMs handle both category-based queries (e.g., “apple”) and attribute-based descriptions (e.g., “red food”) with similar effectiveness.
@@
Temporal tasks pose additional challenges to the existing VLMs, with models consistently failing to identify the correct order of interactions, since it requires multi-hop reasoning. Duration tracking is similarly difficult as it requires inferring elapsed duration based on specific points in the video. In contrast, for the time-based tasks, we observe that baselines perform better, as these tasks only require single frame retrieval based on the timestamps in the question, without additional reasoning.
Poor performance on multi-goal tasks. From fig. 2, we observe that all VLM baselines achieve near-zero performance on most multi-goal tasks. This indicates that VLMs struggle to track and recall multiple object–receptacle interactions. In our multi-goal evaluation, we employ a single-shot prompt, requiring the high-level VLM agent to predict all target frame indices in one response, primarily to avoid repeated inferences over long sequences given the large number of frames. Qualitatively, we find that VLMs often fail to predict even the correct number of targets needed to solve a task. While our trained Qwen model performs somewhat better, reaching around 20% accuracy, it still falls far short of the attainable 99% success rate on this task.
VLMs can identify target receptacles but not precisely localize them. We analyze VLM performance on tasks where the valid entities are receptacles versus where they are objects. As shown in fig. 3(a), the Semantic Coverage-SR (SC-SR) metric (see section B.3) is substantially higher for receptacle-based tasks. This indicates that VLMs often select frames where the receptacle is visible, but the agent remains far from it. In contrast, distance-threshold success (DTG-SR) improves negligibly, suggesting that very few cases exist where the agent is close to a receptacle but not facing it. For smaller objects, no such improvement is observed (see fig. 3(b)), since small objects are rarely visible from longer distances. Taken together, these results highlight a key limitation of current VLMs: they can semantically detect large targets but fail to construct spatially grounded representations necessary for precise localization.
(a) Relaxed metrics: Recep.
(b) Relaxed metrics: Objects.
(c) HL-SR vs HL-SPL.
(d) Hierarchical policy.
Figure 3: (a, b) Relaxed metrics for object and receptacle tasks. Models show substantial gains on receptacle-based tasks under relaxed metrics, whereas tasks targeting small objects exhibit little improvement. (c) Success Rate vs. SPL. For tasks where the VLM may select among multiple correct frames, all models sometimes choose suboptimal ones, reducing the HL-SPL metric. (d) Hierarchical policy on FindingDory. When Qwen is paired with an ImageNav or Mapping agent, overall SR and SPL drop markedly (as reflected in $L$-SR/$L$-SPL).
Policy decomposition causes performance degradation. To evaluate the hierarchical agent introduced in section 4.1, we use Qwen2.5-VL-7B as the high-level VLM module and pair it with either an ImageNav policy or a mapping-based policy as the low-level agent. The low-level policy is triggered only when the high-level VLM correctly identifies a subgoal. As shown in fig. 3(d), the Low-Level-SR (LL-SR) metric drops sharply across all FindingDory tasks for the ImageNav policy. We attribute this decline to a distribution shift in the visual goals provided at evaluation. During training, the policy sees randomly sampled start–end goal poses, whereas VLM-selected goals often correspond to pick–place routines. These frames largely contain background walls or partial receptacle surfaces, offering limited visual cues for navigation. By contrast, the mapping-based navigation policy is more robust, suffering only a 25% relative drop in SR. However, the SPL of both methods remains in the single digits, underscoring the importance of FindingDory in exposing these shortcomings and motivating the development of more reliable memory and navigation strategies.
Low level navigation policy failures. From  Figure 4(b), we observe that the low-level image-goal navigation policy can fail due to invoking the termination (stop) action incorrectly or a navigation routine timeout (max budget of $1000$ steps including the data collection phase). To further understand this behavior, we analyze if the navigation policy actually looks at the target entity when the stop action is invoked. Interestingly, we find that in several instances the target entity is visible from a slightly larger distance threshold (2.0m) from the chosen task success threshold (1.0 m, in case of receptacles). This often happens when VLMs select goal frames in which the target entity is visible in the 2D image but not necessarily closest in 3D space, and thus the goal frame coordinates are slightly further away than the threshold distance.
(a) Comparison across different video lengths.
(b) ImageNav Policy Errors
@@
Leveraging additional frames. We study the effect of input length on high-level performance by evaluating agents on subsampled interaction videos of varying sizes (fig. 4(a)). This allows us to test whether models benefit from longer histories or merely pick up sparse visual cues. Interestingly, frozen VLMs, although capable of processing long contexts, show little to no improvement with more frames. Their performance often degrades at higher frame counts, suggesting difficulty in extracting relevant signals from densely packed, unfiltered inputs.
In contrast, the fine-tuned model (Qwen SFT) achieves clear gains when trained with longer videos. This shows that with appropriate supervision, models can move beyond shallow matching to exploit richer temporal structure in interaction histories. Nonetheless, even with full finetuning on longer (subsampled) videos, a large gap remains relative to maximum achievable performance (see section B.4). This underscores the need for future methods that enable VLMs to extract and store in a compressed manner (i.e. via memory mechanisms) task-relevant information more effectively from long video sequences within the context window.
Method MCQ Accuracy ($\uparrow$) Numerical Accuracy ($\uparrow$)
Zero-Shot 33.69 21.47
Video-R1-Only SFT 33.33 32.67
Video-R1 + FindingDory SFT 35.06 32.35
Video-R1-Only GRPO 33.57 33.9
Video-R1 + FindingDory GRPO 35.42 34.82
Table 3: VSI-Bench results. Models trained with FindingDory show consistent gains.
Cross-Benchmark Results. In this experiment, we explore whether the spatiotemporal memory capabilities enhanced by training on simulator-based FindingDory data can translate to real world settings. To evaluate this, we use VSI-Bench  Yang et al. (2024), a challenging benchmark specifically designed to measure spatiotemporal reasoning in VLMs on egocentric real-world videos. We first augment the Video-R1-CoT-165k dataset introduced in Feng et al. (2025) with $40000$ samples from the FindingDory dataset. For effective co-training, we construct chain-of-thought traces for the entire training split of the FindingDory dataset (see section B.5). We provide specifics about training and evaluation in  appendix F. We then evaluate the trained model on VSI-Bench, showing the results in  table 3. We observe that co-training with FindingDory improves the SFT baseline performance by $1.7\%$ (MCQ accuracy). Furthermore, RL-finetuning with GRPO using combined data from the FindingDory dataset improves performance with a $1.85\%$ increase over the Video-R1-Only RL checkpoint (MCQ accuracy). Overall, the consistent gains obtained from co-training point to a sim-to-real transfer effect, suggesting that FindingDory can augment scarce real-world data for advancing spatiotemporal understanding in VLMs.
Qualitative Examples. We present multiple example responses from various VLMs in  appendix G. Overall we observe that closed models perform better than open-source counterparts on single-goal tasks (see figs. 18, 21 and 22) but all models suffer on multi-goal tasks (see  fig. 23). We also provide example agent video trajectories at https://findingdory-benchmark.github.io/.